# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 5 Thematic analysis
This section explores the diverse applications of GNN models in different groups. As said, the selected journal papers are divided in groups according to the application domain. The goal is to provide an overview of the various empirical studies, the most common approaches, and the results obtained for different applications.
In each subsection of the following thematic analysis, a list of the benchmark models that appear in the selected journal papers will be provided. They will be classified as: mathematical and statistical models coming from Econometrics, traditional machine learning models, and GNN-based models. As for mathematical and statistical models, they generate forecasts by analyzing a limited amount of data, and they usually struggle with cross-sectional large-scale and temporal long-term dependencies. Traditional machine learning models use more complex architectures, mainly neural networks, but they are not necessarily able to capture intricate spatio-temporal dependencies within the data. GNN-based models instead, attempt to represent these complex dependencies using various techniques, resulting in better, though not perfect, accuracy. In addition, information on the most common datasets will be given, along with the best results reported in the literature. All results presented in the tables of this SLR are taken directly from the tables of the selected papers.
The groups of papers are presented in alphabetical order, with the exception of the ”Generic” applications group (which cannot be directly attributed to a specific case study) and the ”Other topics” group (focusing on specific problems of other disciplines). In subsections with a large number of papers and high homogeneity of the datasets, more precise comparisons of the results and in-depth analyses will be conducted.
@@ 5.1 Energy
The first group of papers is ”Energy” (with 13 out of 156 papers), which includes all the studies related to energy systems. In particular, it includes research on electricity load, energy consumption, and power generation forecasting. This is somehow connected to the ”Environment” subsection (Subsec. 5.2). Indeed, green energy sources are associated with environmental phenomena such as wind speed and solar radiation inflow. However, this subsection focuses on energy applications, whereas Subsec. 5.2 refers to the study of environmental data that may or may not be related to energy applications. Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [art_29]. However, wind speed forecasting may also be of interest for other reasons, and thus it is included in the more general Subsec. 5.2.
@@ 5.1.1 Overview
Energy forecasting is crucial for electricity power grid stability and operational efficiency. Accurate load and generation forecasts ensure that supply meets real-time demand, preventing outages and instability, and maintaining a reliable supply of power. In addition, effective forecasting allows for optimal scheduling of power plants and energy resources, reducing operational costs and improving overall system efficiency. The majority of econometric and machine learning models tend to focus on time series from a single site or consider multiple sites without explicitly capturing the relationships between them. GNN models may be an innovative and effective approach to addressing energy related problems.
An overview of the selected papers reveals a noticeable interest in the application of GNNs in the energy field since 2022. The most frequently studied fields include forecasting of wind ([art_20], [art_24], [art_27]) and solar power ([art_19], [art_21], [art_22], [art_26]), which are renewable energy sources characterized by their intermittency. The two most popular journals among the selected papers are Applied Energy by Elsevier, and Frontiers in Energy Research by Frontiers Media SA. The majority of the corresponding authors are affiliated with institutions in China and in the United States.
@@ 5.1.2 Datasets
The study of the selected papers in the ”Energy” group suggests that there are no particularly common or widely recognized benchmark datasets. Indeed, researchers often focus on specific datasets, typically collected within the university campus or in the country where they are located. For the same reason, the links to these datasets are not always available. In Tab. LABEL:t:link_dataset_energy are listed the links to the public datasets mentioned in the papers.
Table 3: List of public datasets in the ”Energy” group and their corresponding links.
@@
UMass Smart* Dataset [art_17] https://traces.cs.umass.edu/index.php/Smart/Smart
Three of the selected papers ([art_20], [art_22], [art_24]) use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S. Department of Energy (https://www.nrel.gov/). Additional datasets were used for exogenous weather variables and satellite imagery (such as Himawari-8, available at https://www.data.jma.go.jp/mscweb/en/himawari89/space_segment/spsg_sample.html in Ref. [art_19]), but they are of course highly dependent on the location of the study. The data granularity ranges from 1 minute to 1 hour, and the forecasting horizon from tens of minutes to days. They depend on the dataset, the domain of application, and the purpose of the forecast. Most datasets require pre-processing, including missing data interpolation, normalization, and satellite image processing.
5.1.3 Proposed models
As for the taxonomy of the proposed GNN models, 8 out of 13 papers use a convolutional GNN, and the other 5 use an attentional GNN. Regarding the graph structure, the papers typically provide only a brief description. For example, it is rarely highlighted whether the graph structure is static or dynamic, and the way in which it is constructed is often described summarily. In some papers, the graph structure reflects the geographic location of the elements (e.g., wind or solar power plants, as in Refs. [art_24] and [art_26]), whereas in others it is based on the correlation or similarity between time series (as in Ref. [art_19]). In some studies, a pre-defined structure is imposed and the weighted adjacency matrix is learned directly from the model, often using the attention mechanism as in Ref. [art_22].
The most common loss function used to train the models is the MSE. Many papers mention the programming languages Python and Matlab, and few of them mention the Python libraries PyTorch [PyTorch] and TensorFlow [2] for the implementation of the GNN model. In this group, only Ref. [art_17] provides the link to the source code of the proposed residential load forecasting with multiple correlation temporal graph neural networks (RLF-MGNN) model (available at https://codeocean.com/capsule/9294192/tree/v1).
@@ 5.1.4 Benchmark models
Tab. LABEL:t:benchmark_models_energy lists the benchmark models used in the ”Energy” group by at least two papers included in this review.
Table 4: List of benchmark models in the ”Energy” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_15], [art_16], [art_22], [art_25]
Classical machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_15], [art_19]
Convolutional neural network (CNN) [art_15], [art_18], [art_21], [art_22]
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_15], [art_17]
@@
Feed-forward neural network (FNN) [art_15], [art_18], [art_19], [art_25], [art_27]
Gated recurrent unit (GRU) [art_15], [art_25]
@@
$k$-nearest neighbors (KNN) [art_18], [art_24]
Long-short term memory network (LSTM) [art_15], [art_16], [art_19], [art_20], [art_23], [art_24]
@@
Support vector regression (SVR) [art_21], [art_22], [art_24], [art_27]
Transformer [attention] [art_16], [art_19]
GNN methods Spatio-temporal autoregressive model (ST-AR) [ST-AR] [art_21], [art_22]
The majority of the papers only use classical machine learning benchmarks. Some papers use simple mathematical and statistical models, such as linear regression (that we incorporated into ARIMA) or historical average. Only few papers use other GNN-based models as benchmarks, and there are no common benchmarks among them. This is due to the fact that this field has not been extensively explored in the past, and only a very limited number of GNN models have been specifically developed for energy forecasting.
@@ 5.1.5 Results
In all the papers, the authors claim that the proposed GNN model outperforms the benchmarks. The most common error metrics are the MAE, the RMSE, and the Normalized Root Mean Square Error (NRMSE).
It is not possible to compare the accuracy of the discussed models because there are no common benchmark datasets. The only two papers that study the same dataset are written by the same authors (Refs. [art_21] and [art_22]), and their most recent paper includes a comparison with the models proposed in the first one.
@@ 5.2 Environment
The group ”Environment” includes a significant number of the selected papers (27 out of 156), which suggests a considerable research interest on the application of GNNs in this field. However, this group is highly fragmented, as it includes a large number of subdomains and applications, all related to the study of environmental data.
@@ 5.2.1 Overview
The group of environmental studies is broad and includes several disciplines such as physics, climatology, oceanography, and atmospheric science. The breadth of the group is reflected in the number of case studies presented in the selected papers, which cover air quality prediction (in terms of PM2.5 concentration), sea temperature, wind speed, and other applications. Forecasting these quantities can be useful for several reasons. For instance, predicting wind speed or rainfall can assist in anticipating some extreme weather phenomena, and is also related to the prediction of green energy outputs. For example, in Refs. [art_29] and [art_34] the authors estimate the wind power forecasting error obtained by using wind speed forecasting models. As another example, predicting sea surface temperature can be useful for weather forecasting, fishing directions, and disaster warnings [art_28]. Moreover, monitoring and forecasting environmental data are crucial for evaluating the impact of human activities on the environment and tracking the progression of climate change.
The methods for forecasting environmental data are typically categorized as numerical models, statistical models, and machine learning methods. Numerical methods are based on atmospheric models, and are designed to quantify the interaction of different atmospheric variables. The accuracy of these models depends on the availability of data, and the mathematical simulations to get the forecasts may take days or even weeks, thus limiting the ability to make good short-term forecasts. Linear statistical methods leverage long-term dependencies in the recorded data by using regression-like models. However, they cannot capture non-linear relationships in the dataset, which limits their power. Machine learning models, on the other hand, are often a generalization of linear statistical methods that can capture more complex and non-linear relationships in the data.
An overview of the selected papers reveals that there has been a rapid increase in interest in the field among the GNN research community, and more than 20 of the selected papers were published in 2023. The most investigated subjects are sea temperature ([art_28], [art_33], [art_35], [art_36], [art_39], [art_41], [art_47]), wind speed ([art_29], [art_31], [art_34], [art_40], [art_44], [art_46]), and PM2.5 concentration forecasting ([art_2], [art_3], [art_4], [art_6], [art_7], [art_8]). All the selected papers are published in different journals, with the exception of the Engineering Applications of Artificial Intelligence by Elsevier, which has published two of the papers. The majority of the corresponding authors are affiliated with institutions in China.
@@ 5.2.2 Datasets
The analysis of the selected papers shows that in this group of papers there is no reference benchmark dataset. The publicly available datasets cited in the papers are listed in Tab. LABEL:t:link_dataset_environment.
Table 5: List of public datasets in the ”Environment” group and their corresponding links.
@@
China Environmental Monitoring Station PM2.5 [art_2], [art_8] https://www.cnemc.cn/
@@
KnowAir [art_6] https://doi.org/10.1145/3397536.3422208 (source paper)
@@
NOAA sea datasets [art_28], [art_36], [art_39], [art_46], [art_47] https://www.ndbc.noaa.gov/
@@
Five of the selected papers ([art_28], [art_36], [art_39], [art_46], [art_47]) use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S. Department of Commerce (US DOC). The US DOC website permits users to download daily, weekly and monthly mean optimum interpolation sea surface temperature (OISST) data from September 1981 for many geographical locations. In over half of the papers, exogenous variables are incorporated into the model. These include weather data (e.g., humidity, rainfall, pressure, temperature, wind) and, in the case of PM2.5 concentration forecasting, other pollutants (e.g., CO, NO2, O3, PM10, SO2).
The data granularity ranges from 1 minute to 1 month, depending on the case study, and the forecasting horizon from 10 minutes to 240 days. Most datasets require some pre-processing, including interpolation of missing data, removal of outliers and normalization.
5.2.3 Proposed models
Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNN (17 out of 27 papers), followed by 6 attentional GNNs and 3 hybrid architectures. The description in one paper (Ref. [art_5]), however, does not specify the classification of its model within the proposed taxonomy. As for the graph structure, not all of the papers describe it with sufficient precision, and some of them are vague about the graph construction and its static or dynamic nature. In almost half of the papers, the edge weights are learned directly from the model, often using the graph attention mechanism. When the graph structure is defined manually, the most popular criterion is the spatial distance between the sites ($d$), even in its exponential ($e^{-d}$) or inverse proportion ($1/d$) formulation. Finally, there are a few papers where more than one graph structure is fused together in order to capture more complex dynamics, as in Ref. [art_2].
Not all papers specify the loss function used to train the models. Among those that do, MAE and MSE are used in roughly equal proportions. Many papers mention the programming language Python and the library PyTorch for the implementation of the models. Few papers also provide a link to the source code for the proposed model, as shown in Tab. LABEL:t:link_codici_environment.
Table 6: List of source codes of the ”Environment” models in the review.

### v2

@@ 5 Thematic analysis
This section explores the applications of the collected GNN models across different thematic groups. The goal is to provide an overview of available empirical studies, the most common approaches, and the results obtained concerning the discussed themes.
In each following subsection, one per theme, a list of the benchmark models that appear in the selected journal papers will be provided. Within each theme, they will be classified as: mathematical and statistical models coming from Econometrics, non-GNN machine learning models, and GNN-based models. Mathematical and statistical models are characterized by the fact that they generate forecasts by analyzing a limited amount of data, and they usually struggle with cross-sectional large-scale and temporal long-term dependencies. Non-GNN machine learning models use more complex and highly nonlinear architectures, mainly neural networks, but they are not necessarily able to capture intricate spatio-temporal dependencies within the data. GNN-based models, conversely, attempt to represent these complex dependencies using various techniques, resulting in better, though not perfect, accuracy. In addition, information on the most common datasets will be given, along with the best results reported in the literature. All results presented in the tables of this SLR are taken directly from the tables of the selected papers, as we did not conduct any further experiments to obtain additional results. Note that in the tables which list benchmark models we have included references to any GNN model found. In the results tables, instead, we have explicitly referenced only those models proposed in the selected papers, in order to highlight them. Further information about the models used as benchmarks can be found either in the benchmark tables or in the papers where they are used as benchmarks.
The thematic groups of papers are presented in alphabetical order, with the exception of the “Generic” applications group (which cannot be directly attributed to a specific case study) and the “Other topics” group (focusing on specific problems of other disciplines). In subsections with a large number of papers and high homogeneity of the datasets, more precise comparisons of the results and in-depth analyses will be conducted.
@@ 5.1 Energy
The first thematic group is “Energy” (with 33 out of 281 papers), which includes all the studies related to energy systems. In particular, it includes research on electricity load, energy consumption, and power generation forecasting. These topics are somehow connected to those of the “Environment” subsection discussed later (Subsec. 5.2). For example, green energy sources are associated with environmental phenomena such as wind speed and solar radiation inflow. Our rationale for this selection is that here we focus on energy applications, whereas Subsec. 5.2 will refer to the study of environmental data that may or may not be related to energy applications. Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [art_29]. However, wind speed forecasting may also be of interest for other reasons, and thus it is included in the more general Subsec. 5.2.
@@ 5.1.1 Overview
Energy forecasting is crucial for electricity power grid stability and operational efficiency. Accurate load and generation forecasts ensure that supply meets real-time demand, preventing outages and instability, and maintaining a reliable supply of power. In addition, effective forecasting allows for optimal scheduling of power plants and energy resources, reducing operational costs and improving overall system efficiency. In the papers within this theme, the majority of econometric and non-GNN machine learning models tend to focus on time series from a single site or consider multiple sites without explicitly capturing the relationships between them. GNN models may be an innovative and effective approach to addressing energy related problems.
An overview of the selected papers reveals a relevant interest in the application of GNNs in the energy field starting from 2022. The most frequently studied fields include forecasting of wind [art_20, art_24, art_27, art_332, art_333] and solar power [art_19, art_21, art_22, art_26, art_349], which are renewable energy sources characterized by their intermittency. The most popular journal among the selected papers is Applied Energy by Elsevier, with 5 published papers. The majority of the corresponding authors are affiliated with institutions in China.
@@ 5.1.2 Datasets
The study of the selected papers in the “Energy” group suggests that there are no particularly common or widely recognized benchmark datasets within this field. Indeed, researchers often focus on specific datasets, typically collected within the university campus or in the country where they are located. For the same reason, the links to these datasets are not always available. In Tab. LABEL:t:link_dataset_energy are listed the links to the public datasets mentioned in the papers.
Table 7: List of public datasets in the “Energy” group and their corresponding links.
@@
Commercial and Residential Load [art_348] https://www.osti.gov/biblio/1788456
@@
Individual Household Electric Power Consumption [art_344] https://archive.ics.uci.edu/ml/datasets/individual+household+electric+power+consumption
@@
PJM Hourly Energy Consumption [art_390] https://www.pjm.com/markets-and-operations/data-dictionary
PVGIS [art_349] http://re.jrc.ec.europa.eu/pvgis/apps4/pvest.php
@@
Residential heat and electricity demand [art_341] https://ieee-dataport.org/open-access/8-years-hourly-heat-and-electricity-demand-residential-building
SGCC dataset [art_342] http://www.sgcc.com.cn/
Southern Power Grid [art_344] https://www.csg.cn/
UMass Smart* Dataset [art_17] https://traces.cs.umass.edu/docs/traces/smartstar/
Three of the selected papers [art_20, art_22, art_24] use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S. Department of Energy (https://www.nrel.gov/). Additional datasets were used for exogenous weather variables and satellite imagery (such as Himawari-8, available at https://www.data.jma.go.jp/mscweb/en/himawari89/space_segment/spsg_sample.html in Ref. [art_19]), but they are of course highly dependent on the location of the study. The data time granularity ranges from 1 minute to 1 day, and the forecasting horizon from tens of minutes to days. These quantities depend on the dataset, the domain of application, and the purpose of the forecast. Most datasets require pre-processing, including missing data interpolation, normalization, and satellite image processing.
5.1.3 Types of models
As for the taxonomy of the proposed GNN models, 17 out of 33 papers use a convolutional GNN, and the other 9 use an attentional GNN. Six papers from 2024 use both convolutional and attentional approaches, and one paper [art_338] uses the original recurrent approach on a graph defined via the visibility algorithm. As to the graph structure used, the papers typically provide only a brief description. For example, only few papers highlight whether the graph structure is static or dynamic, and the way in which it is constructed is often described summarily. In some papers, the graph structure reflects the geographic location of the elements (e.g., wind or solar power plants, as in Refs. [art_24, art_26, art_349]), whereas in others it is based on the correlation or similarity between time series (as in Ref. [art_19]). In some studies, a pre-defined structure is imposed and the weighted adjacency matrix is learned directly from the model, often using the attention mechanism as in Ref. [art_22]. The majority of papers in this group rely on a static graph.
The most common loss function used to train these models is the MSE. Many papers mention the use of the programming languages Python and Matlab, and few of them mention the use of the Python libraries TensorFlow [2], PyTorch [PyTorch], and PyTorch Geometric [PyTorch_Geometric] for the implementation of the GNN model. In this group, only Ref. [art_17] provides the link to the source code of the proposed residential load forecasting with multiple correlation temporal graph neural networks (RLF-MGNN) model (available at https://codeocean.com/capsule/9294192/tree/v1).
@@ 5.1.4 Benchmark models
Tab. LABEL:t:benchmark_models_energy lists the benchmark models used in the “Energy” group by at least two papers included in this review.
Table 8: List of benchmark models in the “Energy” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_15], [art_16], [art_22], [art_25], [art_333], [art_337], [art_338], [art_341], [art_343], [art_349]
Historical average (HA) [art_350], [art_390]
Vector ARIMA [art_332], [art_333], [art_343], [art_350], [art_390]
Non-GNN machine learning methods Autoformer [Autoformer] [art_333], [art_341]
Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_15], [art_19], [art_339]
Convolutional neural network (CNN) [art_15], [art_18], [art_21], [art_22], [art_343], [art_347]
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_15], [art_17], [art_351]
@@
Feed-forward neural network (FNN) [art_15], [art_18], [art_19], [art_25], [art_27], [art_351]
Gated recurrent unit (GRU) [art_15], [art_25], [art_334], [art_338], [art_343], [art_347]
@@
$k$-nearest neighbors (KNN) [art_18], [art_24], [art_342], [art_343]
Long-short term memory network (LSTM) [art_15], [art_16], [art_19], [art_20], [art_23], [art_24], [art_333], [art_334], [art_335], [art_340], [art_341], [art_343], [art_344], [art_347], [art_348], [art_349], [art_353]
Long-short term time series network (LSTNet) [LSTNet] [art_15], [art_332], [art_334]
@@
Seq2Seq [Seq2seq] [art_335], [art_343]
Support vector regression (SVR) [art_21], [art_22], [art_24], [art_27], [art_338], [art_342], [art_343], [art_344], [art_348]
Transformer [attention] [art_16], [art_19], [art_335], [art_343]
Extreme gradient boosting (XGBoost) [art_25], [art_334], [art_335], [art_341]
GNN methods Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_349], [art_350], [art_390]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_332], [art_340], [art_341]
Spatio-temporal autoregressive model (ST-AR) [ST-AR] [art_21], [art_22]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_339], [art_341], [art_348], [art_351], [art_353]
The majority of the papers only use non-GNN machine learning benchmarks. Some papers use simple mathematical and statistical models, such as the ARIMA model and the historical average. Only few papers use GNN-based models as benchmarks, and the most popular benchmark among them is the spatio-temporal graph convolutional network (STGCN). Before 2024, the only benchmark used by two papers was the spatio-temporal autoregressive model (ST-AR). This is due to the fact that this thematic field has not been extensively explored in the past, and only a very limited number of GNN models have been specifically developed for energy forecasting.
@@ 5.1.5 Results
In general, the authors of the papers in this group claim that their proposed GNN model outperforms the benchmarks. However, in some cases, it turns out that basic statistical methods may achieve good results as well [art_390]. The most common error metrics used are the MAE, the RMSE, and the Normalized Root mean squared Error (NRMSE). Some papers calculate these error metrics across different seasons to evaluate the model’s performance under different conditions. An example is provided in Ref. [art_349], where the authors also use the GNNExplainer [interpret_2] to assess the influence of different variables.
It is not possible to compare the accuracy of the models included in this group, because no common benchmark datasets were used. The only two papers that study the same dataset are written by the same authors (Refs. [art_21, art_22]), and their most recent paper includes a comparison with the models proposed in the first one.
@@ 5.2 Environment
The thematic group “Environment” includes a relevant number of selected papers (43 out of 281). This suggests a considerable research interest on the application of GNNs in this field. However, this group is highly fragmented, as it includes a large number of subdomains and applications, all related to the study of environmental data.
@@ 5.2.1 Overview
The group of papers related to environmental topics is broad, and includes papers on disciplines such as physics, climatology, oceanography, and atmospheric science. The breadth of the group is reflected in the number of case studies, which cover air quality prediction (in terms of PM2.5 concentration), sea temperature, wind speed, and other applications. Forecasting these quantities can be useful for several reasons. For instance, predicting wind speed or rainfall can assist in anticipating some extreme weather phenomena, and is also related to the prediction of green energy outputs. For example, in Refs. [art_29] and [art_34] the authors estimate the wind power forecasting error obtained by using wind speed forecasting models. As another example, predicting sea surface temperature can be useful for weather forecasting, fishing directions, and disaster warnings [art_28]. Moreover, monitoring and forecasting environmental data are crucial for evaluating the impact of human activities on the environment and tracking the progression of climate change.
The methods for forecasting environmental data are typically categorized as numerical methods, statistical methods, and machine learning methods. Numerical methods are based on atmospheric models, and are designed to quantify the interaction of different atmospheric variables. The accuracy of these models depends on the availability of data, and the mathematical simulations needed to get the forecasts may take days or even weeks, thus limiting the ability to make good short-term forecasts. Linear statistical methods leverage long-term dependencies in the recorded data by using regression-like models. However, they cannot capture non-linear relationships in the dataset, which limits their power. Machine learning methods, on the other hand, are often advanced generalizations of linear statistical methods that can capture more complex and non-linear relationships in the data.
An overview of the selected papers reveals that there has been a rapid increase in interest on this theme among the GNN research community too, and more than 35 of the selected papers were published from 2023. The most investigated data are related to sea temperature (Refs. [art_28, art_33, art_35, art_36, art_39, art_41, art_47]), wind speed (Refs. [art_29, art_31, art_34, art_40, art_44, art_46, art_366]), and PM2.5 concentration forecasting (Refs. [art_2, art_3, art_4, art_6, art_7, art_8, art_354, art_361, art_371]). Almost every selected paper is published in a different journal, with the exception of Energy, Engineering Applications of Artificial Intelligence, Journal of Cleaner Production, Knowledge-Based Systems, Science of the Total Environment by Elsevier, each with two published papers. The majority of the corresponding authors are affiliated with institutions in China.
@@ 5.2.2 Datasets
The analysis of the selected papers shows that in this thematic group there is no standardized benchmark dataset. The publicly available datasets cited in the papers are listed in Tab. LABEL:t:link_dataset_environment.
Table 9: List of public datasets in the “Environment” group and their corresponding links.
@@
China Environmental Monitoring Station PM2.5 [art_2], [art_8], [art_371] https://www.cnemc.cn/
@@
HadCRUT5 [art_370] https://doi.org/10.1029/2019JD032361 (source paper)
Jinan PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
KnowAir [art_6] https://github.com/shuowang-ai/PM2.5-GNN
@@
NOAA sea datasets [art_28], [art_36], [art_39], [art_46], [art_47], [art_365] https://www.ndbc.noaa.gov/
@@
Overbetuwe groundwater data [art_359] https://github.com/mlttac/GroundwaterFlowGNN
@@
US Geological Survey [art_355] https://earthquake.usgs.gov/earthquakes/search/
Yangtze River PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
Six of the selected papers [art_28, art_36, art_39, art_46, art_47, art_365] use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S. Department of Commerce (US DOC). The US DOC website permits users to download daily, weekly and monthly mean optimum interpolation sea surface temperature (OISST) data from September 1981 for many geographical locations. In over half of the papers, exogenous variables are incorporated into the model. These include weather data (e.g., humidity, rainfall, pressure, temperature, wind) and, in the case of PM2.5 concentration forecasting, concentration of pollutants (like CO, NO2, O3, PM10 and SO2).
Overall, many papers also consider exogenous variables, in particular weather data, as they have the greatest influence on environmental data.
The data time granularity ranges from 1 minute to 1 month, depending on the case study, and the forecasting horizon ranges from 10 minutes to more than one year. Most datasets require some type of pre-processing, including interpolation of missing data, removal of outliers and normalization.
5.2.3 Types of models
Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNNs (28 out of 43 papers), followed by 9 attentional GNNs, and by 3 hybrid architectures. The description in 3 of the papers (Refs. [art_5, art_354, art_361]), however, does not allow us to specify the classification of the model within the proposed taxonomy. As for the graph structure, not all of the papers describe it with sufficient precision, and some of them are vague about the graph construction and its static or dynamic nature. In almost half of the papers, the edge weights are learned directly from the model, often using the graph attention mechanism. When the graph structure is defined manually, the most popular criterion used is the spatial distance between the sites ($d$), even in its exponential ($e^{-d}$) or inverse proportion ($1/d$) formulation, or similarity measures, such as cosine similarity and correlation of the time series. Finally, there are a few papers where more than one graph structure is fused together in order to capture more complex dynamics, as in Ref. [art_2]. Almost all the papers in this thematic group deal with multivariate time series, with the exception of Ref. [art_370], which studies a univariate time series of global average temperature. In this work, the authors treat each time step $t$ as a node $v_{t}$ in the graph, and the adjacency matrix is constructed in such a way that $v_{t}$ is influenced by $v_{t-1}$ and $v_{t+1}$.
Not all papers openly write down the loss function used to train the models. Among those that do, the most popular loss functions are MSE and MAE. Many papers mention the use of the programming language Python and of the library PyTorch for the implementation of the models. Few papers also provide a link to the source code for the proposed model, as shown in Tab. LABEL:t:link_codici_environment.
Table 10: List of source codes of the “Environment” models in the review.